ABSORPTION SPECTOSCOPY of the n = 0 → n = 2 PSEUDOROTATIONAL BAND in TETRAHYDRO-FURAN

\underline{DMITRY G. MELNIK}, SANDHYA GOPALAKRISHNAN, TERRY A. MILLER, The Ohio State University, Dept. of Chemistry, Laser Spectroscopy Facility, 120 W. 18th Avenue, Columbus, Ohio 43210; REBECCA A. H. BUTLER and FRANK C. DE LUCIA, The Ohio State University, Dept. of Physics, Microwave Laboratory, 174 W. 18th Avenue, Columbus OH 43210.

The previously reported^{a} rotational structure observed in the region of 193-335 GHz using the pulsed jet FASSST scan spectrometer in tetrahydrofuran (THF) is assigned as originating from pseudo-rotational bands. A total of 121 lines were observed of which 112 lines were assigned to n = 0 → n = 2 band, and 9 are tentatively assigned to n = 1 → n = 2 band using the molecular constants of THF reported in microwave studies^{b}. All available data from the microwave and the present studies were globally fit to produce an improved set of molecular constants. The origin of the n = 0 → n = 2 is determined to be equal to 269613.25 MHz. An experimentally observed deviation of the selection rules from those previously derived^{c} is discussed.

aD.Melnik, S.Gopalakrishnan, T.A.Miller, and F.C.De Lucia, 54^{th} International Symposium on Molecular Spectroscopy

bR. Meyer, J. C. Lopez, J. L. Alonso, S. Melandri, P. G. Favero, and W. Caminati J. Chem. Phys, 111, 7871 (1999)

cJ. A. Greenhouse and H. L. Strauss, J. Chem. Phys. 50, 124 (1969)